Given the functions: Evaluate the function (g)) for 20. Give your answer in exact form ( (20)) is
Answer #1

Sol) Given, the functions f(x)=-or, Now, f(g(x)) =f(g(20)) g(x) /5x, h(x)=3x-6 x=20) ( : 20 =r(,5(20)) g(20) 5(20) =f(A00) -f